A man got a 10% increase in his salary. If his new salary is ₹1,54,000, find his original salary. 

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

Let the original salary be x.
Given that, the new salary is₹1,54,000.
Original salary + Increment = New salary
Given that the increment is 10% of the original salary.
∴ x+(10/100 × x) =154000
x + x/10=154000
11x/10=154000
x = 154000×10/11
= 140000
∴ The original salary was₹1,40,000

Thanks
Harshit Singh
3 years ago
Dear Student

Let the original salary be x.
Given that, the new salary is₹1,54,000.
Original salary + Increment = New salary
Given that the increment is 10% of the original salary.
∴ x+(10/100 × x) =154000
x + x/10=154000
11x/10=154000
x = 154000×10/11
= 140000
∴ The original salary was₹1,40,000

Thanks